# Large language model psychometrics: A systematic review of evaluation, validation, and enhancement

### v2

@@ 5.4.1 Closed-ended output and scoring
LLM performance on structured tests is typically evaluated using logit-based probabilistic analysis and closed-ended output scoring. In multiple-choice or Likert-scale settings, some studies analyze token-level logits—especially the first token—to infer latent traits, detect response entropy, or compare LLM output distributions against human data [Pellert et al., 2024, Santurkar et al., 2023, Dominguez-Olmedo et al., 2024, Arehalli et al., 2022]. Psycholinguistic tests often use logit-based metrics such as surprisal (the negative log probability of a token), which quantifies predictability and cognitive effort, allowing researchers to align LLM uncertainty patterns with those of humans [Wang et al., 2024b, Steuer et al., 2023].
Conversely, closed-ended outputs, which are explicit numerical scores or categorical selections, can be analyzed using predefined scoring protocols. For Likert-scale responses, scores are typically averaged or aggregated based on established rubrics (e.g., VSM13 [Ye et al., 2025a] and DIT [Khandelwal et al., 2024]). In standardized cognitive tests, model outputs are evaluated against ground-truth labels (e.g., accuracy in arithmetic tasks) or rule-based criteria (e.g., established standards in the Verbal Comprehension Index) [Galatzer-Levy et al., 2024].
@@ 5.4.2 Open-ended output and scoring
Rule-based scoring predominantly relies on the lexical hypothesis [Allport and Odbert, 1936], analyzing the presence and frequency of specific keywords or phrases. For example, Jiang et al. [2024b] apply LIWC (Linguistic Inquiry and Word Count) features [Pennebaker et al., 2001] to evaluate LLM personalities, and Fischer et al. [2023] employ a theory-driven value dictionary [Ponizovskiy et al., 2020] for value measurement. However, such lexicon-based approaches struggle to capture semantic nuance [Ye et al., 2025a]. Some rule-based methods go beyond keyword matching; for instance, Healey et al. [2024] detect bias by checking for deviations from equivalent treatment in responses.
@@
Model-based scoring is prevalent in unstructured testing for its flexibility and scalability. Many studies train models to evaluate LLM responses, such as fine-tuning BERT variants on the MyPersonality dataset for personality scoring [Hilliard et al., 2024]. Others fine-tune LLMs with psychometric inventories or human annotations to classify value valence [Ye et al., 2025a, Yao et al., 2024, 2025b, Sorensen et al., 2024a]. Scoring can be conducted at the item level (e.g., Generative Psychometrics: item parsing, scoring, aggregation [Ye et al., 2025a]) or at the response level, evaluating the entire output [Yao et al., 2024, 2025b]. Some models are tailored to specific frameworks (e.g., Schwartz’s Value Theory [Yao et al., 2024]), while others are trained to be generalists by leveraging broader fine-tuning data and LLM prior knowledge [Ye et al., 2025a]. Model-based evaluators are especially common for personality and value assessment, where textual cues map readily to target dimensions. Many studies now use the LLM-as-a-judge approach [Gu et al., 2024] for scoring open-ended responses. For instance, Li et al. [2024d], Zheng et al. [2025] prompt LLMs to rate responses on personality, ToM, and motivation, validating consistency with other LLMs or human raters. This method is especially useful in unstructured simulation settings, such as SI measurement across goal completion, relationship maintenance, and social rule adherence [Zhou et al., 2024b, Wang et al., 2024d, a, Mou et al., 2024]. Embedding models are also used to measure similarity between LLM outputs and prototypical examples [Huang et al., 2024c, Cahyawijaya et al., 2024, Amirizaniani et al., 2024].
@@
Human scoring is employed when rigorous evaluation, adhering strictly to standard psychometric manuals, is necessary. For instance, Elyoseph et al. [2023] engage psychologists to assess the contextual suitability of LLM responses using the Levels of Emotional Awareness Scale (LEAS). Castello et al. [2024] conduct an examination of cognitive biases in LLMs through human evaluation and linguistic comparison. Healey et al. [2024] present a semi-automated pipeline to classify LLM responses into nuanced bias types. Other examples include ToM evaluation in open-ended responses [Amirizaniani et al., 2024] and cognitive evaluation based on the Montreal Cognitive Assessment (MoCA) [Dayan et al., 2024].
@@ 6 Psychometric validation
Unlike AI benchmarking, which focuses on system performance, psychometrics prioritizes theoretical grounding, standardized protocols, and reproducibility. Psychometric validation ensures that tests are reliable, valid, and fair. As an emerging field, LLM Psychometrics lacks standardization in test design and administration, and recent work is beginning to address these issues. We outline the landscape of psychometric validation in Fig. 4.
@@ 6.1 Reliability and consistency
Reliability, a core aspect of psychometric validation, measures how consistently a test performs—over time (test-retest), across versions (parallel forms), and among evaluators (inter-rater). Researchers have adapted these metrics for LLMs. For example, Li et al. [2024d] propose a benchmark covering five reliability forms: internal consistency, parallel forms, inter-rater, option position robustness, and adversarial attack robustness. Other studies [Zheng et al., 2025, Huang et al., 2023a, Ceron et al., 2024, Shu et al., 2024] examine LLM output consistency across repeated trials, prompt variations, and languages. Notably, some metrics capture both random and systematic errors, so inconsistencies can undermine both reliability and validity.
@@
Some advanced LLMs show stable Big Five personality trait responses across settings [Huang et al., 2023a], with larger and instruction-tuned models achieving even higher reliability using role-playing prompts [Serapio-García et al., 2023]. LLMs also display consistent value orientations across paraphrased, related, and translated questions, as well as different response formats [Moore et al., 2024]. Strong inter-rater reliability is reported when LLMs act as judges, though personality stability and value consistency vary by model and cultural context [Li et al., 2024d]. However, other studies report less favorable results. LLMs are notoriously sensitive to prompt variations: small changes in prompt form [Röttger et al., 2024, Ren et al., 2024], option order [Li et al., 2024d, Dominguez-Olmedo et al., 2024], or syntax [Hu and Levy, 2023] can systematically shift outputs, undermining parallel forms reliability. Studies also find inconsistencies between next-token logits and forced-choice responses in psycholinguistic tests [Hu and Levy, 2023], and between forced-choice and free-form responses in value [Kovač et al., 2024, Ye et al., 2025a] and political opinion surveys [Röttger et al., 2024]. In cognitive tests, minor changes to ToM tasks often cause poor performance, indicating low parallel forms reliability and reliance on fragile heuristics [Ullman, 2023, Shapira et al., 2023a, Holterman and Deemter, 2023]. The Words and Deeds Consistency Test (WDCT) [Xu et al., 2025b] further shows LLMs’ verbal and behavioral responses are inconsistent across domains.
@@
Reliability findings are mixed and depend on factors such as construct, reliability type, test format, prompting, and model. LLMs are generally more reliable in personality assessments [Huang et al., 2023a, Serapio-García et al., 2023] than in political opinion evaluations [Dominguez-Olmedo et al., 2024, Röttger et al., 2024]. Inter-rater agreement varies across instruments and domains [Bodroža et al., 2024]. LLMs often show high internal consistency (Cronbach’s alpha $>0.8$) in closed-choice personality tests [Huang et al., 2023a, Serapio-García et al., 2023, Zheng et al., 2025], but low parallel form reliability due to prompt sensitivity [Li et al., 2024d, Shu et al., 2024, Gupta et al., 2024]. Consistency improves in structured tests with strong prompt control [Rozen et al., 2024, Serapio-García et al., 2023, Klinkert et al., 2024], but declines with less standardized prompts [Petrov et al., 2024] or controversial topics [Moore et al., 2024]. Reliability also varies by model: base models are more consistent than fine-tuned ones on value-laden questions [Moore et al., 2024], and higher model safety correlates with greater value consistency [Ye et al., 2025b]. Advanced models also perform better in role-playing personality tests, increasing reliability [Serapio-García et al., 2023].
@@ 6.2.1 Content validity
Content validity requires that a test fully represents the intended construct. Key challenges include data contamination from established tests and insufficiently vetted novel items. Directly applying human tests to LLMs risks contamination, as LLMs may have seen test items or similar content during training [Hagendorff et al., 2024, Jiang et al., 2024c]. This can inflate performance or bias trait measurement. To mitigate this, researchers reformat tests or generate new items; see § 5.2. On the other hand, reformatted or newly generated items may inadequately capture the construct or introduce extraneous factors. This may be due to a less rigorous crowd-sourcing process for benchmark curation, or the inherent biases and limited capabilities of the AI models used for stimulus synthesis. Evaluating content validity for custom-curated and model-generated items is crucial but rarely conducted in LLM Psychometrics.
@@ 6.2.2 Construct validity
Construct validity determines if a test truly measures its intended construct. Alaa et al. [2025] highlight its critical yet often lacking role in medical LLM benchmarks. Broader literature suggests that construct validity is challenged by LLMs’ unique internal abstractions, systematic response patterns, and social desirability bias.
Unique internal abstractions
@@
LLMs and humans may fundamentally differ in their internal representations of psychological constructs. Kovač et al. [2023] argue that LLMs display contextually adaptive traits, unlike the stable characteristics seen in humans. Sühr et al. [2023] show that LLM personality types do not align with the clear separation of Big Five traits found in humans. Similarly, Peereboom et al. [2024] find that human-oriented HEXACO tests are inapplicable to LLMs, as key personality factors may be absent. For values, Ye et al. [2025a] report that LLM self-reports do not fit Schwartz’s model, and while open-ended responses improve alignment, it remains an imperfect fit; Hadar-Shoval et al. [2024] also find LLMs encode values distinct from humans. Overall, researchers agree that robust LLM psychometric evaluation requires new operational definitions, distinct from or only loosely analogous to human traits.
@@
In this regard, pioneering research presents value systems tailored to LLMs’ unique internal abstractions [Ye et al., 2025b, Biedma et al., 2024]. Ye et al. [2025b] introduce a generative psycho-lexical approach based on GPV value measurement [Ye et al., 2025a], demonstrating through various tasks that their value factors are valid and outperform Schwartz’s system. The distinct nature of LLM constructs requires rigorously designed tests: Ma et al. [2025] present the Core Sentiment Inventory (CSI) for measuring LLM emotional tendencies and personality; Fang et al. [2024] create a psychologically grounded mathematics benchmark to improve validity, model item difficulty, and enable human comparison; and Lee et al. [2024d] propose the TRAIT test, achieving higher validity and reliability than other personality measures.
@@
Response set refers to systematic answer patterns regardless of item content. Some LLMs show option position bias, favoring responses labeled "A" in political opinion tests [Dominguez-Olmedo et al., 2024, Li et al., 2024d]. After correcting for this bias, Dominguez-Olmedo et al. [2024] find LLM responses to political questions are uniformly random, rendering demographic alignment analyses meaningless. Similar findings are observed in LLM personality [Song et al., 2023, Sühr et al., 2023] and value tests [Ye et al., 2025a]. Sühr et al. [2023] also note LLMs often affirm both the original and reverse-coded items, while Ye et al. [2025a] observe some LLMs consistently prefer low or high ratings. These observations suggest self-report tests may be unsuitable for LLM Psychometrics.
@@
Social desirability bias—the tendency to give socially acceptable rather than genuine responses—is observed in LLMs as in humans. Salecha et al. [2024] find LLMs skew toward favorable traits in personality tests, inflating extraversion and lowering neuroticism scores. Similarly, Ye et al. [2025b] show LLMs avoid reporting less desirable values (e.g., hedonism) in direct self-reports [Biedma et al., 2024], but reveal them in indirect, contextual prompts. While tuning for social desirability is important, overemphasis risks producing homogeneous models that lack diverse perspectives and fail to meet varied user needs.
@@
Cross-lingual psychometric tests reveal notable inconsistencies in LLM measurement results across languages [Romero et al., 2024]. Despite multilingual capabilities, LLMs often lack consistent psychological trait assessments between languages [Cahyawijaya et al., 2024]. These inconsistencies complicate validation, requiring careful separation of model biases, translation issues, and cultural differences represented by the models.

### v3

@@ 5.4.1 Closed-ended output and scoring
LLM performance on structured tests is typically evaluated using logit-based probabilistic analysis and closed-ended output scoring. In multiple-choice or Likert-scale settings, some studies analyze token-level logits—especially the first token—to infer latent traits, detect response entropy, or compare LLM output distributions against human data [pellert2024ai_psychometrics, santurkar2023whose, dominguez2024questioning, pl6_arehalli2022syntactic]. Psycholinguistic tests often use logit-based metrics such as surprisal (the negative log probability of a token), which quantifies predictability and cognitive effort, allowing researchers to align LLM uncertainty patterns with those of humans [pl3_wang2024how, pl10_steuer2023large].
Conversely, closed-ended outputs, which are explicit numerical scores or categorical selections, can be analyzed using predefined scoring protocols. For Likert-scale responses, scores are typically averaged or aggregated based on established rubrics (e.g., VSM13 [ye2025gpv] and DIT [khandelwal2024moral]). In standardized cognitive tests, model outputs are evaluated against ground-truth labels (e.g., accuracy in arithmetic tasks) or rule-based criteria (e.g., established standards in the Verbal Comprehension Index) [l1_galatzer2024cognitive].
@@ 5.4.2 Open-ended output and scoring
Rule-based scoring predominantly relies on the lexical hypothesis [allport1936trait], analyzing the presence and frequency of specific keywords or phrases. For example, jiang2024personallm apply LIWC (Linguistic Inquiry and Word Count) features [pennebaker2001linguistic] to evaluate LLM personalities, and fischer2023does employ a theory-driven value dictionary [ponizovskiy2020development] for value measurement. However, such lexicon-based approaches struggle to capture semantic nuance [ye2025gpv]. Some rule-based methods go beyond keyword matching; for instance, healey2024evaluating detect bias by checking for deviations from equivalent treatment in responses.
@@
Model-based scoring is prevalent in unstructured testing for its flexibility and scalability. Many studies train models to evaluate LLM responses, such as fine-tuning BERT variants on the MyPersonality dataset for personality scoring [hilliard2024eliciting_personality]. Others fine-tune LLMs with psychometric inventories or human annotations to classify value valence [ye2025gpv, yao2024value_fulcra, yao2025clave, sorensen2024value]. Scoring can be conducted at the item level (e.g., Generative Psychometrics: item parsing, scoring, aggregation [ye2025gpv]) or at the response level, evaluating the entire output [yao2024value_fulcra, yao2025clave]. Some models are tailored to specific frameworks (e.g., Schwartz’s Value Theory [yao2024value_fulcra]), while others are trained to be generalists by leveraging broader fine-tuning data and LLM prior knowledge [ye2025gpv]. Model-based evaluators are especially common for personality and value assessment, where textual cues map readily to target dimensions. Many studies now use the LLM-as-a-judge approach [gu2024llm-as-a-judge] for scoring open-ended responses. For instance, li2024quantifying_ai_psychology, zheng2025lmlpa prompt LLMs to rate responses on personality, ToM, and motivation, validating consistency with other LLMs or human raters. This method is especially useful in unstructured simulation settings, such as SI measurement across goal completion, relationship maintenance, and social rule adherence [si3_zhou2023sotopia, si4_wang2024sotopia, si6_wang2024towards, si8_mou2024agentsense]. Embedding models are also used to measure similarity between LLM outputs and prototypical examples [huang2024designing, cahyawijaya2024high, tom17_amirizaniani2024llms].
@@
Human scoring is employed when rigorous evaluation, adhering strictly to standard psychometric manuals, is necessary. For instance, elyoseph2023chatgpt engage psychologists to assess the contextual suitability of LLM responses using the Levels of Emotional Awareness Scale (LEAS). castello2024examining conduct an examination of cognitive biases in LLMs through human evaluation and linguistic comparison. healey2024evaluating present a semi-automated pipeline to classify LLM responses into nuanced bias types. Other examples include ToM evaluation in open-ended responses [tom17_amirizaniani2024llms] and cognitive evaluation based on the Montreal Cognitive Assessment (MoCA) [l3_dayan2024age].
@@ 6 Psychometric validation
Unlike AI benchmarking, which focuses on system performance, psychometrics prioritizes theoretical grounding, standardized protocols, and reproducibility. Psychometric validation ensures that tests are reliable, valid, and fair. As an emerging field, LLM Psychometrics lacks standardization in test design and administration, and recent work is beginning to address these issues. We outline the landscape of psychometric validation in Fig.˜4. Model fairness is also paramount; however, given the extensive literature on AI fairness, we refer readers to dedicated surveys [mehrabi2021survey, gallegos2024bias_fairness_survey].
@@ 6.1 Reliability and consistency
Reliability, a core aspect of psychometric validation, measures how consistently a test performs—over time (test-retest), across versions (parallel forms), and among evaluators (inter-rater). However, applying these concepts to LLMs requires a fundamental shift in perspective compared to human psychology. In human psychometrics, the subject (the person) is assumed to have relatively stable traits, and reliability metrics (e.g., Cronbach’s $\alpha$) primarily evaluate the quality of the instrument. In contrast, LLM Psychometrics faces a dual challenge: the instability of the instrument (due to flawed design) and the stochastic nature of the subject (due to models’ context sensitivity).
@@
Researchers adapt these metrics for LLMs to tackle the dual challenges. For example, li2024quantifying_ai_psychology propose a benchmark covering five reliability forms: internal consistency, parallel forms, inter-rater, option position robustness, and adversarial attack robustness. Other studies [zheng2025lmlpa, huang2023revisiting_the_reliability, ceron2024beyond, shu2024you_dont_need] examine LLM output consistency across repeated trials, prompt variations, and languages.
When measuring reliability, many studies implicitly or explicitly adopt the "LLM as a population" perspective when applying reliability metrics [serapio2023personality_traits_in_llms]. Unlike human psychometrics, where reliability usually assesses consistency across a population of distinct individuals, here reliability often verifies whether a single tool consistently measures a construct within a single LLM, which is treated as a population of diverse generated responses elicited by diverse prompts. The prompts can be diversified in terms of model profiling, item rephrasing, or both. If a model’s responses remain internally consistent across prompt variations (e.g., high Cronbach’s $\alpha$), it demonstrates that the test reliably captures the latent traits of that specific model-population.
Other works treat the LLM as a single individual, focusing on the stability of its outputs for a specific test. When treating LLM as a single entity, reliability is assessed through metrics like parallel forms reliability, option position robustness, and adversarial attack robustness, which measure whether the model’s responses remain stable despite minor variations in the test items or presentation format.
@@
Some advanced LLMs show stable Big Five personality trait responses across settings [huang2023revisiting_the_reliability], with larger and instruction-tuned models achieving even higher reliability using role-playing prompts [serapio2023personality_traits_in_llms]. LLMs also display consistent value orientations across paraphrased, related, and translated questions, as well as different response formats [moore2024large]. Strong inter-rater reliability is reported when LLMs act as judges, though personality stability and value consistency vary by model and cultural context [li2024quantifying_ai_psychology]. However, other studies report less favorable results. LLMs are notoriously sensitive to prompt variations: small changes in prompt form [rottger2024political, ren2024valuebench], option order [li2024quantifying_ai_psychology, dominguez2024questioning], or syntax [pl7_hu2023prompt] can systematically shift outputs, undermining parallel forms reliability. Studies also find inconsistencies between next-token logits and forced-choice responses in psycholinguistic tests [pl7_hu2023prompt], and between forced-choice and free-form responses in value [kovavc2024stick, ye2025gpv] and political opinion surveys [rottger2024political]. In cognitive tests, minor changes to ToM tasks often cause poor performance, indicating low parallel forms reliability and reliance on fragile heuristics [tom1_ullman2023large, tom4_shapira2023clever, tom13_holterman2023does]. The Words and Deeds Consistency Test (WDCT) [xu2025large] further shows LLMs’ verbal and behavioral responses are inconsistent across domains.
@@
Reliability findings are mixed and depend on factors such as construct, reliability type, test format, prompting, and model. LLMs are generally more reliable in personality assessments [huang2023revisiting_the_reliability, serapio2023personality_traits_in_llms] than in political opinion evaluations [dominguez2024questioning, rottger2024political]. Inter-rater agreement varies across instruments and domains [bodroza2024personality_testing_of_llms]. LLMs often show high internal consistency (Cronbach’s alpha >0.8>0.8) in closed-choice personality tests [huang2023revisiting_the_reliability, serapio2023personality_traits_in_llms, zheng2025lmlpa], but low parallel form reliability due to prompt sensitivity [li2024quantifying_ai_psychology, shu2024you_dont_need, gupta2024self_assessment_tests]. Consistency improves in structured tests with strong prompt control [rozen2024llms, serapio2023personality_traits_in_llms, klinkert2024driving_generative_agents], but declines with less standardized prompts [petrov2024limited_ability_of_llms] or controversial topics [moore2024large]. Reliability also varies by model: base models are more consistent than fine-tuned ones on value-laden questions [moore2024large], and higher model safety correlates with greater trait consistency [ye2025generative_psycho_lexical, tosato2025persistent]. Advanced models also perform better in role-playing personality tests, increasing reliability [serapio2023personality_traits_in_llms].
@@ 6.2.1 Content validity
Content validity requires that a test fully represents the intended construct. Key challenges include data contamination from established tests and insufficiently vetted novel items. Directly applying human tests to LLMs risks contamination, as LLMs may have seen test items or similar content during training [hagendorff2024machine, jiang2024investigating]. This can inflate performance or bias trait measurement. To mitigate this, researchers reformat tests or generate new items; see §˜5.2. On the other hand, reformatted or newly generated items may inadequately capture the construct or introduce extraneous factors. This may be due to a less rigorous crowd-sourcing process for benchmark curation, or the inherent biases and limited capabilities of the AI models used for stimulus synthesis. Evaluating content validity for custom-curated and model-generated items is crucial but rarely conducted in LLM Psychometrics.
@@ 6.2.2 Construct validity
Construct validity determines if a test truly measures its intended construct. Literature suggests that construct validity is challenged by LLMs’ reasoning shortcuts, unique internal abstractions, systematic response patterns, and social desirability bias.
Procedural validity.
Procedural validity refers to the extent to which the internal processes producing a response align with the theoretical construct being measured. Standard psychometrics often relies on a functionalist assumption: correct answers imply underlying competence. However, it has been argued that for LLMs, high accuracy is a poor proxy for intelligence if the process of reasoning is flawed or based on “shortcuts” [mitchell2025artificial, stevenson2024can, lewis2024evaluatingrobustness]. Research warns against the “anthropomorphic fallacy”, which assumes that human-like outputs evidence human-like cognition [Lin2025fallacies]. Without verifying procedural validity, LLM Psychometrics risks conflating sophisticated statistical mimicry with genuine understanding. Pioneering works exemplified by beger2025ai evaluate LLMs’ abstract reasoning capabilities across modalities, analyzing both output accuracy and generated rules. The results indicate top reasoning models like o3 achieve human-level accuracy but frequently rely on unintended "shortcut" rules. han2025dual further delve into the internal mechanisms of LLMs and reveal that LLMs express values through two distinct pathways: intrinsic expression, reflecting values learned during training, and prompted expression, elicited by specific instructions. These mechanisms rely on partially distinct internal components (e.g., value vectors and neurons), demonstrating that similar outputs may arise from different processes. Further work must move beyond output accuracy to scrutinize the reasoning process itself.
Construct equivalence.
A foundational concern in LLM Psychometrics is construct equivalence: whether a psychological construct such as personality or values carries the same meaning when applied to LLMs as it does for humans. This problem is structurally analogous to the “imposed etic” issue in cross-cultural psychometrics, where instruments developed in one cultural context are applied to another without first verifying conceptual equivalence [berry1969cross_cultural]. Transposing this logic to LLM Psychometrics: applying human personality inventories to LLMs presupposes that the constructs are equivalent across the human-LLM divide, which is an assumption that requires explicit empirical testing rather than being taken for granted [vandenberg2000review].
When investigating the latent psychological structure of LLMs, researchers usually treat a single LLM or multiple LLMs with diverse prompts as a population. This approach ensures sufficient variance for factor analysis or other structural equation modeling techniques to identify underlying dimensions. Based on this statistical probing, recent studies find that LLMs and humans may fundamentally differ in their internal representations of psychological constructs.
@@
kovac2023llms_as_superpositions argue that LLMs display contextually adaptive traits, unlike the stable characteristics seen in humans. suhr2023challenging show that LLM personality types do not align with the clear separation of Big Five traits found in humans. Similarly, peereboom2024cognitive find that human-oriented HEXACO tests are inapplicable to LLMs, as key personality factors may be absent. For values, ye2025gpv report that LLM self-reports do not fit Schwartz’s model, and while open-ended responses improve alignment, it remains an imperfect fit; hadar2024assessing also find LLMs encode values distinct from humans. Taken together, these findings constitute evidence of construct non-equivalence: the same instrument does not measure the same latent property across the human-LLM divide. Researchers therefore agree that robust LLM psychometric evaluation requires new operational definitions, distinct from or only loosely analogous to human traits.
@@
To address this, the field is moving toward AI-native psychometrics. Pioneering research presents value systems tailored to LLMs’ unique latent abstractions [ye2025generative_psycho_lexical, biedma2024beyond_human_norms]. ye2025generative_psycho_lexical introduce a generative psycho-lexical approach based on GPV value measurement [ye2025gpv], demonstrating through various tasks that their value factors are valid and outperform Schwartz’s system. The distinct nature of LLM constructs requires rigorously designed tests: ma2025leveraging present the Core Sentiment Inventory (CSI) for measuring LLM emotional tendencies and personality; fang2024patch create a psychologically grounded mathematics benchmark to improve validity, model item difficulty, and enable human comparison; and lee2024do_llms_have_distinct propose the TRAIT test, achieving higher validity and reliability than other personality measures.
@@
Response set refers to systematic answer patterns regardless of item content. Some LLMs show option position bias, favoring responses labeled "A" in political opinion tests [dominguez2024questioning, li2024quantifying_ai_psychology]. After correcting for this bias, dominguez2024questioning find LLM responses to political questions are uniformly random, rendering demographic alignment analyses meaningless. Similar findings are observed in LLM personality [song2023have_large_language, suhr2023challenging] and value tests [ye2025gpv]. suhr2023challenging also note LLMs often affirm both the original and reverse-coded items, while ye2025gpv observe some LLMs consistently prefer low or high ratings. These observations suggest self-report tests may be unsuitable for LLM Psychometrics.
@@
Social desirability bias—the tendency to give socially acceptable rather than genuine responses—is observed in LLMs as in humans. salecha2024large find LLMs skew toward favorable traits in personality tests, inflating extraversion and lowering neuroticism scores. Similarly, ye2025generative_psycho_lexical show LLMs avoid reporting less desirable values (e.g., hedonism) in direct self-reports [biedma2024beyond_human_norms], but reveal them in indirect, contextual prompts. While tuning for social desirability is important, overemphasis risks producing homogeneous models that lack diverse perspectives and fail to meet varied user needs.
@@
Cross-lingual psychometric tests reveal notable inconsistencies in LLM measurement results across languages [romero2024do_gpt_language_models]. Despite multilingual capabilities, LLMs often lack consistent psychological trait assessments between languages [cahyawijaya2024high]. These inconsistencies complicate validation, requiring careful separation of model biases, translation issues, and cultural differences represented by the models.